# Laminar chaos in systems with random and chaotically time-varying delay

### v1

@@ V.1 Test for laminar chaos
These structures are general in the sense that they are visible as long as the fluctuations of the delay $\tau(t)$ and therefore the fluctuations of the lengths of the state intervals are not too large such that the average roundtrip time $T$ and therefore the offset of the subspaces is finite, as well as the fluctuations around the subspace are small enough such that the clusters for different values of $p$ are distinguishable. Therefore they can be used to extract the sequence $p_{n}$ from the time series by dividing the $d$-dimensional return map of the durations into such point clusters. After that, the value $p_{n}$ of the $n$th laminar phase can be determined simply by looking to which cluster the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ belongs. One possible method for solving this classification problem can be derived from the k-subspace clustering method introduced in [130]. The points of the $d$-dimensional return map are classified by the Euclidean distance to the subspaces defined by Eq. (25) with $T_{l}=T$. In detail, $p_{n}$ is set of the value $p$ such that the distance between the corresponding subspace and the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ is minimal. Since $T$ is unknown, we scan for the value of $T$ that maximizes the nonlinear correlations of the return map given by the points $(z_{n},z_{n+p_{n}})$ quantified by $(1-Q)$. Such a scan computed from our exemplary time series is shown in Fig. 12, where the algorithm was applied to a $4$-dimensional return map of durations such that $4$ clusters corresponding to $p=1,2,3,4$ were identified. The cluster classification obtained at the value $T$ that maximizes $(1-Q)$ is shown in Fig. 13. It nicely coincides with the result shown in Fig. 11(b), which was obtained using the knowledge of the time-varying delay, except that the estimate $T\approx 2.78$ differs from the expected value $T=3.0$ leading to the visible shift of the clusters relative to the subspaces. Extracting the sequence $p_{n}$ and plotting $(z_{n},z_{n+p_{n}})$ reveals the nonlinearity $f$ of the delay system as shown in Fig. 14. However, the nonlinear correlations are slightly lower than the values obtained when using the knowledge of the time-varying delay, compare Fig. 10. This is caused by the partial overlap of the point clusters for larger values of $p$, for instance, close to the point $(\delta_{n},\delta_{n+1},\delta_{n+2})=(0,T,0)$, where the subspace for $p=2$ (line) intersects the subspace for $p=3$ (plane). Also the points in between the dense point clouds lead to wrong values $p_{n}$ since by the algorithm they are assigned to the clusters however they do not meet the assumption that $p$ must be equal for all laminar phases contributing to the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ to be close to one of the affine subspaces. This probably can be improved by updating the proof-of-concept algorithm used here, for instance, by using duration return maps with different dimensions $d$. Then larger values of $p_{n}$ can be identified in higher-dimensions, where lower dimensions are more suitable for lower values of $p$ since $p_{n}$ has to be constant for at least $d$ laminar phases to lie close to one of the subspaces, which becomes less likely if $d$ is increased.
@@ V.2 Identification of pseudolaminar chaos
A millimeter-sized drop of silicone oil can walk horizontally while bouncing vertically on a vertically vibrating bath of the same liquid [131, 132]. The droplet on each bounce generates localized standing waves that decay slowly in time, and the droplet interacts with these self-generated waves on subsequent bounces to propel itself horizontally. We call such a walking droplet an active wave-particle entity and it has the following three key features: (i) the droplet and the wave co-exist as a wave-particle entity; the repeated bouncing of the droplet sustain the damped waves which in turn guide the walking motion of the droplet, (ii) the system is active in the sense of an active particle [133] since it absorbs energy from the vibrating bath and converts it into self-propulsion, and (iii) the system is non-Markovian and has path memory since the waves generated by the droplet decay very slowly in time, so the droplet is not only influenced from its most recent wave, but also from the waves generated in the distant past. By averaging over the fast time scale of periodic vertical bouncing, one obtains an integro-differential trajectory equation for horizontal walking motion [134]. This trajectory equation for walking motion in $1$D in dimensionless form for a particle generating idealized cosine waves and confined to a harmonic potential takes the form of Lorenz-like nonlinear ordinary differential equations (ODEs) as follows [135, 134, 136, 137, 138]:
@@
Figure 15: Mechanism behind pseudo-laminar chaos for the system in Eq. (26). A laminar phase persists as long as the chaotically varying parameter approximately fulfills $r(t)<1$, corresponding to a stationary state of the dynamical system, and a deviation from the laminar phases is triggered when we have $r(t)>1$, corresponding to the instability of the stationary state. (a) Time series of the variable $x$ in the system of Lorenz-like ODEs in Eq. (26) with $r=r(t)$. (b) Shifted chaotically varying parameter $r(t)-1$. See also Supplemental Video 1.
@@
Figure 16: Quantity $(1-Q)$ as a function of the parameter $T$ of the clustering algorithm for pseudolaminar chaos, where the $(d=4)$-dimensional return map of the durations of the (pseudo)laminar phases was analyzed. The return maps of the levels and durations of the laminar phases for the parameter $T\approx 68$ (dashed line) that maximizes $(1-Q)$ are used to classify the type of chaotic dynamics.
@@
We now apply our test for laminar chaos introduced in the previous section to the time series $x(t)$ with the length of $10^{6}$ time units  [^8] The time series was generated via the MATLAB function ode45 while setting the options ’RelTol’ and ’AbsTol’ both to $10^{-10}$. , where the first $1000$ time units are dropped to let the transients relax. We use the exact same methods to identify the laminar phases and to extract the durations $\delta_{n}$ and the levels $x_{n}$ of the nearly constant phases. Using the $(d=4)$-dimensional return map of the durations $\delta_{n}$, we try to identify the sequence of the $p_{n}$ by identifying the clusters that are associated to the specific values. Therefore we vary the parameter $T$ that determines the location of the subspaces defining the shape of the clusters, compute the sequence $p_{n}$ from the results and compute the value of the quantity $(1-Q)$ for the resulting return map given by the points $(x_{n},x_{n+p_{n}})$. As shown in Fig. 16, the nonlinear correlations quantified by $(1-Q)$ are maximized for $T=68$, where a step size of $\Delta T=1$ was used. The three-dimensional return map of the durations together with the subspaces for $p=1,2,3$ is shown in Fig. 17. The majority of points were assigned to the point cluster with $p=1$, where the remaining ones make of $0.6\%$ and belong to exceptional long nearly constant phases mostly exceeding the plot range of Fig. 17. This means that the return map of the levels $x_{n}$ given by the points $(x_{n},x_{n+p_{n}})$ is dominated by the relation between two subsequent nearly constant phases. The quantities $(1-Q)$ and MIC take values around $0.5$ and therefore indicate some sort of correlations, which are caused by the harmonic potential pulling the particle position towards $x=0$. However, the return map as shown in Fig. 18 is far from resembling a one-dimensional map, which is contradictory to laminar chaos and is expected since this type of dynamics is generated by a completely different mechanism. So the updated test for laminar chaos enables us to distinguish this type of pseudolaminar chaotic dynamics from laminar chaos if we do not only rely on the values of quantities like $(1-Q)$ and MIC but also take into account the actual structure of the return map of the levels.

### v2

@@ V.1 Test for laminar chaos
If the system and therefore the time varying delay $\tau(t)$ is unknown, the sequence $p_{n}$ in principle can be reconstructed from the durations $\delta_{n}$ of the laminar phases using the $d$-dimensional return maps given by the points $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$. In Fig. 12, the two and three-dimensional return maps of the durations $\delta_{n}$ of the laminar phases are shown, which were computed by subtracting subsequent burst positions obtained from our exemplary time series. The color coding of the values of the $p_{n}$ from our preceding analysis reveals characteristic structures that can be exploited for obtaining the $p_{n}$ even when we do not know anything about the system. The structures arise from this trivial fact: The sum of the durations of the laminar phases inside a solution segment are equal to the length of its domain, which is the state interval whose length equals the delay. As the delay is bounded this leads to clustering around (p1)(p-1)-dimensional affine subspaces (see Appendix A for details), which are illustrated in black in Fig. 12(b). Via the proof-of-concept clustering algorithm provided in Appendix A, these clusters can be identified and the value of $p_{n}$ is assigned by looking to which cluster the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ belongs. Plotting $(z_{n},z_{n+p_{n}})$ finally reveals the nonlinearity $f$ of the delay system as shown in Fig. 11. However, the nonlinear correlations are slightly lower than the values obtained when using the knowledge of the time-varying delay, compare Fig. 10. This is caused by the partial overlap of the point clusters for larger values of $p$, for instance, close to the point $(\delta_{n},\delta_{n+1},\delta_{n+2})=(0,T,0)$, where the subspace for $p=2$ (line) intersects the subspace for $p=3$ (plane). Also the points in between the dense point clouds lead to wrong values $p_{n}$ since by the algorithm they are assigned to the clusters however they do not meet the assumption that $p$ must be equal for all laminar phases contributing to the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ to be close to one of the affine subspaces. This probably can be improved by updating the proof-of-concept algorithm used here, for instance, by using duration return maps with different dimensions $d$. Then larger values of $p_{n}$ can be identified in higher-dimensions, where lower dimensions are more suitable for lower values of $p$ since $p_{n}$ has to be constant for at least $d$ laminar phases to lie close to one of the subspaces, which becomes less likely if $d$ is increased.
@@ V.2 Identification of pseudolaminar chaos
A millimeter-sized drop of silicone oil can walk horizontally while bouncing vertically on a vertically vibrating bath of the same liquid [132, 133]. The droplet on each bounce generates localized standing waves that decay slowly in time, and the droplet interacts with these self-generated waves on subsequent bounces to propel itself horizontally. We call such a walking droplet an active wave-particle entity and it has the following three key features: (i) the droplet and the wave co-exist as a wave-particle entity; the repeated bouncing of the droplet sustains the damped waves which in turn guide the walking motion of the droplet, (ii) the system is active in the sense of an active particle [134] since it absorbs energy from the vibrating bath and converts it into self-propulsion, and (iii) the system is non-Markovian and has path memory since the waves generated by the droplet decay very slowly in time, so the droplet is not only influenced from its most recent wave, but also from the waves generated in the distant past. By averaging over the fast time scale of periodic vertical bouncing, one obtains an integro-differential trajectory equation for horizontal walking motion [135]. This trajectory equation for walking motion in 11D in dimensionless form for a particle generating idealized cosine waves and confined to a harmonic potential takes the form of Lorenz-like nonlinear ordinary differential equations (ODEs) as follows [136, 135, 137, 138, 139]:
@@
Figure 13: Mechanism behind pseudo-laminar chaos for the system in Eq. (25). A laminar phase persists as long as the chaotically varying parameter approximately fulfills $r(t)<1$, corresponding to a stationary state of the dynamical system, and a deviation from the laminar phases is triggered when we have $r(t)>1$, corresponding to the instability of the stationary state. (a) Time series of the variable $x$ in the system of Lorenz-like ODEs in Eq. (25) with $r=r(t)$. (b) Shifted chaotically varying parameter r(t)1r(t)-1. See also Supplemental Video 1.
@@
We now apply our test for laminar chaos introduced in the previous section to an exemplary time series with the length of $10^{6}$ time units  [^10] The time series was generated via the MATLAB function ode45 while setting the options ’RelTol’ and ’AbsTol’ both to $10^{-10}$., where the first $1000$ time units are dropped to let the transients relax. Further technical details can be found in Appendix A. We find that the quantities (1Q)(1-Q) and MIC take values around $0.5$ and therefore indicate some sort of correlations, which are caused by the harmonic potential pulling the particle position towards $x=0$. However, the return map as shown in Fig. 14 is far from resembling a one-dimensional map, which is contradictory to laminar chaos and is expected since this type of dynamics is generated by a completely different mechanism. So the updated test for laminar chaos enables us to distinguish this type of pseudolaminar chaotic dynamics from laminar chaos if we do not only rely on the values of quantities like (1Q)(1-Q) and MIC but also take into account the actual structure of the return map of the levels.
@@ Acknowledgements.
Appendix A Clusters of the durations of laminar phases and their detection
Figure 15: Quantity (1Q)(1-Q) as a function of the parameter $T$ of the clustering algorithm. For each $T$, the sequence $p_{n}$ is extracted from the $(d=4)$-dimensional durations return map. Then (1Q)(1-Q) is computed for the points $(x_{n},x_{n+p_{n}})$. The maximum at $T=2.78$ (dashed line) is located within the range of the time-varying delay (dotted lines).
Figure 17: Quantity (1Q)(1-Q) as a function of the parameter $T$ of the clustering algorithm for pseudolaminar chaos, where the $(d=4)$-dimensional return map of the durations of the (pseudo)laminar phases was analyzed. The return maps of the levels and durations of the laminar phases for the parameter $T\approx 68$ (dashed line) that maximizes (1Q)(1-Q) are used to classify the type of chaotic dynamics.
In the following we provide the technical details of the generalization of the toolbox for the detection of laminar chaos to system with random and chaotically time-varying delay. A distinctive feature of laminar chaos is given by the fact that the levels $z_{n}$ are governed by the dynamics of the one-dimensional iterated map $z_{n+p}=f(z_{n})$, where $p$ is the number $p$ of laminar phases per state interval. Therefore the reconstruction of this map from a time series of an unknown system is crucial for the detection of laminar chaos. While $p$ is constant for periodic and quasiperiodic delays and therefore easy to find, for random and chaotically time-varying delays it is time dependent, $p=p_{n}$, rendering trial-and-error approaches impossible. In the first part we present a systematic method how the number $p_{n}$ of laminar phases per state interval can be extracted from return maps of the durations $\delta_{n}$ of the laminar phases of a time series. This method is benchmarked in the second part using time series that show laminar chaos and pseudolaminar chaos, where the assigned sequences of $p_{n}$ lead to return maps $z_{n+p}=f(z_{n})$ that allow a clear distinction of these type of dynamics without knowing the system as demonstrated in Sec. V
For the reconstruction of the number $p_{n}$ of laminar phases per state interval, we extract the sequence of durations $\delta_{n}$ of laminar phases from a time series, where $n$ is the number of the laminar phase in the time series. We then exploit the geometrical structures that appear in $d$-dimensional return maps constructed from points $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$. The structures arise from this trivial fact: Given that a solution segment $z_{k}(t)$ contains only complete laminar phases, then the sum of the durations of these laminar phases are equal to the length of its domain, which is the state interval $\mathcal{I}_{k}=(t_{k}-\tau(t_{k}),t_{k}]$ and is of one delay length $\tau(t_{k})$. If there are $p$ laminar phases inside the solution segments, it follows that the components of each point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ of our $d$-dimensional return map fulfill (dp+1)(d-p+1) equations given by
$T_{l}=\sum_{m=l}^{l+(p-1)}\delta_{n+m},\quad l=0,1,\dots,d-p.$ (26)
So each $T_{l}$ is simply the sum of $p$ subsequent durations starting with $\delta_{n+l}$. Considering $T_{l}$ as constants and the durations $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ as variables, Eqs. (26) define a (p1)(p-1)-dimensional affine subspace of $\mathbb{R}^{d}$, where the $T_{l}$ define its offset from the origin. Since we assumed that $p$ is equal for all laminar phases contributing to the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$, each $T_{l}$ is a length of a state interval and therefore is equal to a value of the time-varying delay $\tau(t)$. This means that each vector composed of $d$ subsequent durations belongs to an (p1)(p-1)-dimensional affine subspace, where the offset of the subspace fluctuates according to the time-varying delay $\tau(t)$. Averaging these fluctuations by replacing $T_{l}$ with the average roundtrip time of the feedback loop $T=T_{\omega}$ from Eq. (22) with $L\to\infty$ leads to the subspaces represented by the black dot, line, and plane in Fig. 12(b), where the clusters with the corresponding values of $p=1,2,3$ scatter around. For our chaotically time-varying delay with the parameters used here, we obtained an average roundtrip time of $T=T_{\omega}=3.0.$ The loose points away from these clusters belong to the case where $p$ is not equal for all laminar phases contributing to the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$, which happens at a transition from one value of $p_{n}$ to another, see Fig. 9.
These structures are general in the sense that they are visible as long as the fluctuations of the delay $\tau(t)$ and therefore the fluctuations of the lengths of the state intervals are not too large such that the average roundtrip time $T$ and therefore the offset of the subspaces is finite, as well as the fluctuations around the subspace are small enough such that the clusters for different values of $p$ are distinguishable. Therefore they can be used to extract the sequence $p_{n}$ from the time series by dividing the $d$-dimensional return map of the durations into such point clusters. After that, the value $p_{n}$ of the $n$th laminar phase can be determined simply by looking to which cluster the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ belongs.
One possible method for solving this classification problem can be derived from the k-subspace clustering method introduced in [141]. The points of the $d$-dimensional return map are classified by the Euclidean distance to the subspaces defined by Eq. (26) with $T_{l}=T$. In detail, $p_{n}$ is set of the value $p$ such that the distance between the corresponding subspace and the point $(\delta_{n},\delta_{n+1},\dots,\delta_{n+(d-1)})$ is minimal. Since $T$ is unknown, we scan for the value of $T$ that maximizes the nonlinear correlations of the return map given by the points $(z_{n},z_{n+p_{n}})$ quantified by (1Q)(1-Q).
Such a scan computed from a time series of our exemplary laminar chaotic system considered in Sec. V.1 is shown in Fig. 15, where the algorithm was applied to a $4$-dimensional return map of durations such that $4$ clusters corresponding to $p=1,2,3,4$ were identified. The cluster classification obtained at the value $T$ that maximizes (1Q)(1-Q) is shown in Fig. 16. It nicely coincides with the result shown in Fig. 12(b), which was obtained using the knowledge of the time-varying delay, except that the estimate $T\approx 2.78$ differs from the expected value $T=3.0$ leading to the visible shift of the clusters relative to the subspaces.
Now we apply the exact same method to the pseudolaminar chaotic time series $x(t)$ analyzed in Sec. V.2. We identify the laminar phases and extract the durations $\delta_{n}$ and the levels $x_{n}$ of the nearly constant phases. Using the $(d=4)$-dimensional return map of the durations $\delta_{n}$, we try to identify the sequence of the $p_{n}$ by identifying the clusters that are associated to the specific values. Therefore we vary the parameter $T$ that determines the location of the subspaces defining the shape of the clusters, compute the sequence $p_{n}$ from the results and compute the value of the quantity (1Q)(1-Q) for the resulting return map given by the points $(x_{n},x_{n+p_{n}})$. As shown in Fig. 17, the nonlinear correlations quantified by (1Q)(1-Q) are maximized for $T=68$, where a step size of $\Delta T=1$ was used. The three-dimensional return map of the durations together with the subspaces for $p=1,2,3$ is shown in Fig. 18. The majority of points were assigned to the point cluster with $p=1$, where the remaining ones make of $0.6\%$ and belong to exceptional long nearly constant phases mostly exceeding the plot range of Fig. 18. This means that the return map of the levels $x_{n}$ given by the points $(x_{n},x_{n+p_{n}})$ is dominated by the relation between two subsequent nearly constant phases.